33.In the following cash flow stream, X is an unknown cash flow amount (thus, you have an unknown cash flow at the end of year 3 and at the end of year 4). The missing cash flow is the same for both years (that is, X in year 3 is the same dollar amount as X in year 4). If the future value at the end of year 10 of these cash flows deposited into an account that pays 5.95 percent nominal interest per year is $6,120, what is the value of the missing cash flows in years 3 and 4 (that is, what is $X)?

End of Year

Cash Flow

1

$700

2

$700

3

$X

4

$X

5

$700

6

$700
Answer #1

FV of all Cash flows = 6120
700*(1+5.95%)5 + 700*(1+5.95%)4 + X*(1+5.95%)3 + X*(1+5.95%)2 + 700*(1+5.95%)1​​​​​​​ + 700​​ = 6120
3258.2683 + 2.31187 X = 6120
X = (6120 - 3258.2683)/2.31187 = 1237.84

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well